OpenStudy (anonymous):

Term: 9x = 12 x/6 = 8 x/2 = 20 8x = 12 ---------------- Definition: 4x = 6 x/8 = 5 x/12 = 4 3x = 4 ----- your suppose to match and I need help. Can you also explain if you help me?

OpenStudy (notsmartatall):

solve for x

OpenStudy (anonymous):

@notsmartatall yea I know but like ive been learning this stuff in school but never like this so this is hard for me. I dont understand I need someone to help

OpenStudy (anonymous):

@hartnn @Hero PLease help me

OpenStudy (bibby):

Do you understand how to solve for a specific variable? for example \[ \frac{ x }{ 6 } = 8\] how would we isolate the x?

OpenStudy (anonymous):

@bibby by doing its oppisite operation which is multiplying

OpenStudy (bibby):

Exactly. You'd do something similar for each of these. If we multiply we get x = 6 * 8 = 48 Now can you get x/12 = 4?

OpenStudy (anonymous):

@bibby 48 because I multiplied

OpenStudy (bibby):

Congrats. We found a match. 6 more equations and we've got ourselves a correct answer

OpenStudy (anonymous):

@bibby yea but I have to match it which the other equation which I dont understand how to do

OpenStudy (bibby):

All you can really do at this point (I think) is solve each one and then match them after the solving is done. For example, x/2=20 and x/8 = 5. You won't know they're matches until you solve but after that you can say they match

OpenStudy (anonymous):

@bibby OH! I get it so if the othe equation equals the same thing they match.

OpenStudy (bibby):

Exactly :) It's just pattern matching. Do you want to go through the problems and have me make sure it's right?